# T-3DGS: Removing Transient Objects for 3D Scene Reconstruction

### v1

@@ 3.2.1 Feature Extraction
For each training iteration, we extract DINO v2 features [23] from both the input image $I$ and corresponding rendering $\hat{I}$ producing feature maps $f,\hat{f}$ respectively. We choose DINO v2 for several key reasons: (1) its self-supervised training enables robust semantic understanding without class-specific biases, (2) it demonstrates strong performance in distinguishing object boundaries and semantic regions even for previously unseen objects, (3) compared to alternatives like DIFT [40] features, DINO v2 offers significantly faster computation times, making it more practical for iterative training processes. These features provide a rich foundation for our TMP to make accurate decisions about scene dynamics without explicit supervision.
3.2.2 Transient Probability Masks
The extracted features are processed through our TMP module, which implements logistic regression to compute per-feature probability masks: $P^{f}=\sigma(\mathbf{W}\mathbf{f})$, where $\sigma(\cdot)$ is the sigmoid function, $\mathbf{W}$ represents the TMP weights. Since the features are patch-wise ($14\times 14$ px), we upsample predicted probability masks $P^{f}$ and $\hat{P}^{f}$ to the original image resolution using bilinear interpolation, obtaining pixel-wise transient masks $P$ and $\hat{P}$ respectively. Notably, TMP learns to distinguish transient objects without explicit supervision, guided only by the optimization objectives described below.
3.2.3 Loss Function
The TMP is optimized using three complementary losses:
$\mathcal{L}_{\text{transient}}=\mathcal{L}_{\text{RGB}}+\lambda_{prior}\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{consist}},$ (1)
The RGB loss identifies regions where the rendered image differs from the ground truth: $\mathcal{L}_{\text{RGB}}=\left|\left|(1-P)\odot(I-\hat{I})\right|\right|_{1}$. where $\|\cdot\|_{1}$ denotes the $L_{1}$ norm, and $\odot$ represents the element-wise (Hadamard) product. This loss exploits the observation that dynamic objects typically create larger discrepancies between rendered and ground truth images. By weighting these differences with $(1-P)$, we encourage the classifier to assign higher transient probabilities to regions showing significant reconstruction errors.
The regularization term promotes sparse predictions: $\mathcal{L}_{\text{reg}}=\left|\left|P\right|\right|_{1}$. This sparsity constraint helps produce compact, focused masks while reducing false positives.
Previous works  [8, 49] used a sum of $\mathcal{L}_{\text{RGB}}$ and $\lambda_{prior}\mathcal{L}_{\text{reg}}$ without the consistency loss, which has an explicit solution. For each pixel $(i,j)$, the optimal solution has the form:
$P^{*}_{i,j}=\begin{cases}1&\text{if }|I_{i,j}-\hat{I}_{i,j}|>\lambda_{prior}\\ 0&\text{if }|I_{i,j}-\hat{I}_{i,j}|\leq\lambda_{prior}\end{cases}$ (2)
@@ 3.2.3 Loss Function
However, this methodology exhibits significant limitations when reconstructing geometrically complex static structures. In such cases, even static objects may generate RGB errors exceeding $\lambda_{prior}$ due to reconstruction challenges, leading to their erroneous classification as dynamic elements and subsequent masking. This misclassification ultimately results in degraded reconstruction quality of the static scene components. To address these limitations, additional regularization constraints are necessary.
Our novel consistency term prevents misclassification in challenging regions:
$\mathcal{L}_{\text{consist}}=\left|\left|P\odot\hat{P}\right|\right|_{1}.$ (3)
This term addresses cases where static objects with complex surfaces might show high reconstruction error while being semantically consistent. The underlying principle is that while geometric complexity may lead to reconstruction errors in RGB space, the semantic features of static objects remain consistent between input and rendered images. By penalizing overlap between input and rendered image masks ($P$ and $\hat{P}$), we encourage the detector to focus on genuine semantic differences rather than reconstruction artifacts.
Fig. 3 demonstrates the effectiveness of this approach: after 7000 iterations, the model with consistency loss produces precise masks for transient objects in the input image while generating minimal masks for the rendered image. In contrast, without the consistency term, the TMP exhibits numerous false positives in both input and rendered images, often misclassifying poorly reconstructed static regions as transient objects. This behavior validates that the consistency loss effectively distinguishes between genuine dynamic objects and reconstruction artifacts based on their semantic representations.
Figure 3: The comparison of the TMP predictions with and without $\mathcal{L}_{\text{consist}}$ term. The quality of the predicted transient masks for the input image with consistency loss term is significantly better than without this term. The predicted masks by TMP are shown in green color.
@@ 3.2.4 Training Stability
To ensure stable training, we implement several key strategies. First, we detach $\hat{P}$ from the computational graph during consistency loss computation to prevent the model from learning to distinguish between ground truth and rendered images.
During 3DGS optimization, there are periods when renders may be unreliable: at the beginning of training and after each opacity reset. Following  [15], we address this by implementing two key strategies. First, we delay the start of TMP training until the 3DGS optimization has completed its first 500 iterations, ensuring the initial scene reconstruction has reached sufficient quality. Second, after each opacity reset, we temporarily pause the TMP optimization for 250 iterations while keep training 3DGS, allowing the reconstruction to stabilize before resuming transient detection.
Prior to bilinear interpolation, we dilate the masks by one pixel. This dilation step serves two purposes: it compensates for the coarse feature resolution of the foundation model (which processes images in patches) and helps address imprecise boundary predictions.
These modifications ensure robust training and accurate transient object detection across diverse dynamic scenes.
@@ 3.3 Transient Mask Refinement
Our transient area detection pipeline is robust for current benchmarks, where transient object is always dynamic and changes it’s position from frame to frame. But for semi-transient objects, which may do not change position for some frames it fails and masks only parts of video when it’s dynamic. To address these issue, we introduce a masks propagation process that refines transient masks into temporally consistent, accurate masks with high-resolution boundaries across the entire video sequence through refinement and propagation. Each segmentation consists of a binary mask defining the object’s spatial extent and a unique label that remains consistent throughout the video sequence.
@@ 3.3.1 Spatial Refinement
We employ the Segment Anything Model (SAM) [14] to refine our transient maps $P_{i}$ into more precise masks $M^{\prime}_{i}$. For each connected component $C_{i}^{k}$ in $P_{i}$, we sample up to ten points as prompts for SAM, leveraging its ability to generate high-quality segmentations from sparse inputs to extract a set of object-aware masks $M^{\prime}_{i}=\{M^{\prime}{{}_{i}^{j}}\}_{j=1}^{L_{i}}$, where $L_{i}$ is the number of predicted masks for image $I_{i}$. Due to potential inaccuracies in the boundaries of our masks, some sampled points might occasionally fall on the background rather than the object itself (e.g., a point sampled between the legs of a person). To address this issue, we filter the predicted masks based on their local coverage score:
$CS_{\text{local},i}=\frac{|P_{i}\cap M^{\prime}{{}_{i}^{j}}|}{|M^{\prime}{{}_{i}^{j}}|}$ (4)
@@ 3.3.2 Temporal Refinement
To address potential false negatives, we propagate the refined masks $\{M^{\prime}_{i}\}_{i=1}^{N}$ throughout the video using SAM2 [29] to obtain more consistent masks $\{M_{i}\}_{i=1}^{N}$. Our propagation process consists of three stages:
@@
    Forward Propagation: Iterating from the first to the last frame, to propagate segmentation masks forward.
@@
    Backward Propagation: Iterating Last-to-first frame, to propagate information from future frames backwards.
@@
    Final Propagation: Final Propagation: A final first-to-last pass, considering both past and future frames as context, which helps resolve any temporal inconsistencies.
To manage computational resources efficiently, we introduce a memory size parameter $N_{m}$, which limits the number of frames considered during propagation. At each step, we only maintain and use the segmentations from the $N_{m}$ most nearest frames, balancing between temporal consistency and memory constraints.
Throughout propagation, we manage mask intersections to ensure coherent segmentation. For any pair of masks $M_{i}^{l}$ and $M_{i}^{m}$ with $IoU(M_{i}^{l},M_{i}^{m})>\lambda_{merge}$, we merge them into a single mask, assigning the lower of the two original labels to maintain consistency.
@@ 3.3.3 Dynamic Object Filtration
To filter out false positive transients and ensure robust detection, we introduce a Stability Ratio (SR) metric that combines both spatial overlap accuracy and temporal consistency. For each detected object, the SR is calculated as $SR=\frac{1}{N}\sum_{i=1}^{N}(R_{i}\cdot CS_{\text{global},i})$, where $N$ is the number of valid frames, $R_{i}$ is the mean value of the absolute difference between ground truth and rendered images within the masked region in frame $i$, and $CS_{\text{global},i}=|P_{i}\cap M_{i}|/|M_{\text{max}}|$ is the global coverage score. Here, $P_{i}$ represents the prompt mask in frame $i$, $M_{i}$ is the segmentation mask, and $M_{\text{max}}$ is the maximum size of the object mask across all frames. This global score helps evaluate the object’s consistency relative to its largest observed size. A frame is considered valid and contributes to the SR calculation only if its local coverage score (Eq. 4) exceeds the validation threshold $\lambda_{\text{cov}}^{\text{val}}$. Objects with SR below a threshold $\lambda_{SR}$ are filtered out as potential false detections. This dual coverage score system ensures that objects maintain both spatial accuracy through local coverage and temporal consistency through global coverage and difference image values.
@@ 3.4 Artifact-Free Reconstruction
3DGS tends to generate floating artifacts (”floaters”) near the camera, particularly in challenging regions like those identified by transient masks. These artifacts can saturate gradients and degrade overall reconstruction quality. We address this issue through depth-aware regularization.
We render depth D for each pixel using alpha compositing similar to color rendering: $D=\sum_{i=1}^{M}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value of the i-th Gaussian, $T_{i}$ is the accumulated transmittance, and $\alpha_{i}$ is the opacity value. To suppress floating artifacts while preserving sharp depth discontinuities at object boundaries, we apply anisotropic total variation (TV) regularization to the rendered depth maps: $\mathcal{L}_{\text{depth}}=\text{mean}(|\nabla_{x}D|)+\text{mean}(|\nabla_{y}D|)$, where p indexes over pixels, $\nabla_{x}$ and $\nabla_{y}$ are spatial gradients in x and y directions respectively.
@@ 3.5 Masked Gaussian Splatting Optimization
$M_{i}(x,y)=\begin{cases}1&\text{if (x,y) is in an occluded area}\\ 0&\text{if (x,y) is in a static area},\end{cases}$ (5)
where (x,y) represents pixel coordinates in the image. We apply binary dilation to $M_{i}$ for $N_{e}$ iterations, yielding $M^{*}_{i}$. This operation creates a buffer zone around detected dynamic objects, enhancing the robustness of our static scene reconstruction. The final loss for 3DGS is:
@@
where $I_{i}$, $\hat{I_{i}}$, $\mathcal{L}_{depth}$, $\odot$ and $\|\cdot\|_{1}$ are the same as were defined previously, $L_{SSIM}$ is a structural similarity loss, $\overline{M}^{*}_{i}$ is a negation of $M^{*}_{i}$ that represents static background and $\lambda_{\text{SSIM}}$, $\lambda_{\text{L1}}$ and $\lambda_{\text{depth}}$ are weighting factors.
This formulation allows the model to focus on static scene elements, effectively handling dynamic objects in the reconstruction process. By integrating these steps, our method robustly reconstructs static scenes from unconstrained videos, effectively handling transient distractors.

### v2

@@ 3.2.1 Feature Extraction
For each training iteration, we extract DINOv2 features [16] from both the input image $I$ and the corresponding rendering $\hat{I}$, producing feature maps $f,\hat{f}$ respectively. We choose DINOv2 for several key reasons: (1) its self-supervised training enables robust semantic understanding without class-specific biases, (2) it demonstrates strong performance in distinguishing object boundaries and semantic regions even for previously unseen objects, (3) compared to alternatives like DIFT [27] features, DINOv2 offers significantly faster computation times, making it more practical for iterative training processes. These features serve as a robust foundation, enabling RUP to make accurate decisions about scene dynamics without explicit supervision.
3.2.2 Transient 2D Uncertainty Modeling
As previously discussed, most methods detect transient objects by utilizing reconstruction errors. For example, NeRF On-the-go [20] considers RGB residuals:
$R=||\hat{I}-I||_{2}.$ (1)
It assumes that the residuals follow a normal distribution:
$p(R|\sigma)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{R^{2}}{2\sigma^{2}}\right).$ (2)
Therefore, we can obtain negative log likelihood:
$\mathcal{L}_{u}=\frac{R^{2}}{2\sigma^{2}}+\log\sigma+\frac{\log 2\pi}{2}.$ (3)
Although the approach is reasonable, RGB residuals lack robustness. In particular, high-frequency objects often result in high reconstruction errors, producing incorrect misclassification. Similarly, dynamic objects with colors similar to the background may be classified as static. While DSSIM or DINOv2 cosine distance can mitigate some errors, they introduce their own limitations. DSSIM residuals are susceptible to similar errors as RGB residuals, though to a lesser degree. The DINOv2 cosine distance, while highly robust, suffers from low resolution. Upsampling models, such as FeatUP [6], can address this issue, though they introduce upsampling artifacts.
This motivates a new multivariate formulation of uncertainty modeling. Let the residual be a 2-dimensional vector:
$R=\begin{bmatrix}R_{1}\\ R_{2}\end{bmatrix},$ (4)
where $R_{1}$ and $R_{2}$ correspond to different similarity metrics: 1) DINOv2 cosine distance defined like in WildGaussians, except we upscale it with FeatUP [6] and 2) DSSIM. We consider a multivariate normal distribution with zero mean and covariance matrix $\Sigma$:
$p(R)=\frac{1}{(2\pi)\sqrt{|\Sigma|}}\exp\left(-\frac{1}{2}R^{T}\Sigma^{-1}R\right),$ (5)
where the covariance matrix is given by
$\Sigma=\begin{bmatrix}\sigma_{1}^{2}&\rho\sigma_{1}\sigma_{2}\\ \rho\sigma_{1}\sigma_{2}&\sigma_{2}^{2}\end{bmatrix}.$ (6)
The negative log-likelihood function becomes:
$\mathcal{L}_{u}=-\log p(R)=\frac{1}{2}R^{T}\Sigma^{-1}R+\frac{1}{2}\log|\Sigma|+\log 2\pi.$ (7)
In contrast to previous works  [14, 20, 12] we predict three parameters — $\sigma_{1}$, $\sigma_{2}$, $\rho$ instead of a single $\sigma$. This allows us to combine information about both residuals.
There is a problem in this derivation due to the assumption that residuals to be strictly positive. This implies that our distribution represents only the positive quadrant of the bivariate normal distribution. While this is relevant for the subsequent derivations, it does not affect the likelihood, as it only introduces a scaling factor, which can be ignored during optimization. Interestingly, this is not important for the one-dimensional case, where likelihood depends only on the absolute value of the residual.
We train a neural network that takes DINOv2 features from a reference image as input and makes a per pixel prediction of $\Sigma$, and use our likelihood term in Eq. 7 as a loss function. It should be noted that $\Sigma$ can be noninvertible when $\sigma_{i}=0$ or $\rho=\pm 1$. Although we predict $\sigma_{i}$ using a softplus nonlinearity, and $\rho$ using a tanh nonlinearity to avoid undesirable values, in practice this can lead to 1) numerical instabilities and 2) undesirable values due to the discrete representation of numerical values. The second problem is easy to solve with clamping, in our experience, the first problem was mostly solved by introducing normalization layers into the architecture of the neural network.
3.2.3 Model Architecture
Given that our training objective is considerably more challenging than the one-dimensional modeling of WildGaussians [12], our model requires a larger architecture. However, this also offers an advantage over previous methods, as we can use simple upscale layers to make our prediction denser without sacrificing local/nonlocal smoothing. The details of architecture are provided in the Supplementary Material.
3.2.4 Binary Mask
One approach to obtaining a binary mask using the modeled uncertainty is to set a threshold on one of the predicted values or define a new criterion:
$M=\mathbb{I}(f(\sigma_{1},\sigma_{2},\rho)>C),$ (8)
where $\mathbb{I}$ is the indicator function, $f(\sigma_{1},\sigma_{2},\rho)$ is a chosen criterion and $C$ is a threshold chosen as a hyperparameter.
@@ 3.2.4 Binary Mask
However, this methodology has notable limitations when applied to the reconstruction of geometrically complex static structures. In such cases, even static objects may produce substantial residuals, leading to their misclassification as dynamic elements and their subsequent masking. This misclassification ultimately degrades the reconstruction quality of static scene components. To address these limitations, we introduce necessary regularization constraints.
We note that, even though we train our RUP only on a reference images, we can also obtain a per pixel uncertainty prediction $\hat{\Sigma}$ using an image reconstructed by Gaussian Splatting model. Because our model relies on semantic information of DINOv2 features, we should expect it to make a similar prediction in the static areas and a different one in areas corresponding to the dynamic objects. To estimate this discrepancy, we calculate the Kullback-Leibler (KL) divergence $D_{KL}(\mathcal{N}(0,\Sigma)||\mathcal{N}(0,\hat{\Sigma}))$, which takes following form for two normal distributions:
$D_{KL}(\mathcal{N}(0,\Sigma)||\mathcal{N}(0,\hat{\Sigma}))=\frac{1}{2}(tr(\hat{\Sigma}^{-1}\Sigma)-ln(\frac{|\Sigma|}{|\hat{\Sigma}|})-2).$ (9)
Fig. 3 illustrates how this approach reduces false classifications in static regions. Unlike previous methods that obtain masks by estimating uncertainty, we instead utilize divergence. This allows us to incorporate additional information to enhance the consistency of our masks. Hence, our binary masks are obtained based on the new criterion:
$M=\mathbb{I}(D_{KL}>C).$ (10)
Figure 3: During the initial stages of reconstruction, RUP predicts high uncertainty in challenging regions such as backgrounds or high frequency details. However, since RUP relies exclusively on semantic information, calculating the divergence between reference uncertainty $\Sigma$ and reconstructed uncertainty $\hat{\Sigma}$ effectively suppresses these artifacts. Areas with divergence values above the threshold are highlighted in red, while the final predicted transient mask by RUP is shown in green.
Notably, formula (9) is not exact since we are working with a folded multivariate distribution. We can add an ad hoc assumption that we observe only the absolute value of the residuals and its sign is random. Nevertheless, it has proven to be a useful heuristic. We leave a more careful derivation for future work.
@@ 3.2.5 Training Stability
During 3DGS optimization, there are periods when renders may be unreliable, particularly at the beginning of training and after each opacity reset. Building on  [12], we address this by implementing two key strategies. First, we delay the start of RUP training until the 3DGS optimization has completed its first 500 iterations, ensuring the initial scene reconstruction has reached sufficient quality. Second, after each opacity reset, we temporarily pause the RUP optimization for 250 iterations while keep training 3DGS, allowing the reconstruction to stabilize before resuming transient detection. We also use scheduled sampling technique from SpotLessSplats [22].
3.2.6 Mask Dilation
We also dilate our masks, depending on the resolution of the scene. This dilation step serves multiple purposes, primarily covering shadows and reflections caused by objects. These modifications ensure robust training and accurate detection of transient objects in diverse dynamic scenes.
@@ 3.3 Transient Mask Refinement
Our transient area detection pipeline is robust for current benchmarks, where transient objects are always dynamic and change their positions from frame to frame. However, for semi-transient objects, which may not change positions for some frames, it fails and masks only parts of the video when they are dynamic. To address this issue, we introduce a mask propagation process that refines transient masks into temporally consistent, accurate masks with high-resolution boundaries across the entire video sequence through refinement and propagation. Each segmentation consists of a binary mask that defines the object’s spatial extent and a unique label, consistent throughout the video sequence.
@@ 3.3.1 Spatial Refinement
We use the Segment Anything Model (SAM) [11] to refine our transient maps, $P_{i}$, into more precise masks, $M^{\prime}_{i}$. For each connected component $C_{i}^{k}$ in $P_{i}$, we sample up to ten points as prompts for SAM, leveraging its ability to generate high-quality segmentations from sparse inputs to extract a set of object-aware masks $M^{\prime}_{i}=\{M^{\prime}{{}_{i}^{j}}\}_{j=1}^{L_{i}}$, where $L_{i}$ is the number of predicted masks for image $I_{i}$. Due to potential inaccuracies in the boundaries of our masks, some sampled points might occasionally fall on the background rather than the object itself (e.g., a point sampled between the legs of a person). To address this, we filter the predicted masks based on their local coverage score:
$CS_{\text{local},i}=\frac{|P_{i}\cap M^{\prime}{{}_{i}^{j}}|}{|M^{\prime}{{}_{i}^{j}}|}.$ (11)
@@ 3.3.2 Temporal Refinement
To address potential false negatives, we propagate the refined masks, $\{M^{\prime}_{i}\}_{i=1}^{N}$, throughout the video using SAM2 [19] to obtain more consistent masks, $\{M_{i}\}_{i=1}^{N}$. Our propagation process consists of three stages:
@@
    Forward Propagation: Iterating from the first frame to the last, propagating the segmentation masks forward.
@@
    Backward Propagation: Iterating from the last frame to the first, propagating information from future frames backward.
@@
    Final Propagation: A final first-to-last pass, considering both past and future frames as context, which helps to resolve temporal inconsistencies.
To manage computational resources efficiently, we introduce a memory size parameter, $N_{m}$, which limits the number of frames considered during propagation. At each step, we maintain and use segmentations from $N_{m}$ nearest frames, balancing temporal consistency with memory constraints.
During propagation, we manage mask intersections to ensure consistent segmentation. For any pair of masks $M_{i}^{l}$ and $M_{i}^{m}$ where $IoU(M_{i}^{l},M_{i}^{m})>\lambda_{merge}$, we merge them into a single mask, assigning the lower of the two original labels to maintain consistency.
@@ 3.3.3 Dynamic Object Filtration
To filter out false positive transients and ensure robust detection, we introduce the Stability Ratio (SR) metric, which combines spatial overlap accuracy and temporal consistency. For each detected object, the SR is calculated as $SR=\frac{1}{N}\sum_{i=1}^{N}(R_{i}\cdot CS_{\text{global},i})$, where $N$ is the number of valid frames, $R_{i}$ is the mean value of the absolute difference between ground truth and rendered images within the masked region in frame $i$, and $CS_{\text{global},i}=|P_{i}\cap M_{i}|/|M_{\text{max}}|$ is the global coverage score. Here, $P_{i}$ represents the prompt mask in frame $i$, $M_{i}$ is the segmentation mask, and $M_{\text{max}}$ is the maximum size of the object mask across all frames. This global score evaluates the object’s consistency relative to its largest observed size. A frame is considered valid and contributes to the SR calculation only if its local coverage score (Eq. 11) exceeds the validation threshold $\lambda_{\text{cov}}^{\text{val}}$. Objects with $SR$ below a threshold $\lambda_{SR}$ are filtered out as potential false detections. This dual coverage score system ensures that objects maintain both spatial accuracy through local coverage and temporal consistency through global coverage and difference image values.
@@ 3.4 Artifact-Free Reconstruction
3DGS tends to generate floating artifacts (”floaters”) near the camera, particularly in challenging regions like those identified by transient masks. These artifacts can saturate gradients, thereby degrading overall reconstruction quality. We address this issue through depth-aware regularization.
We render the depth $D$ for each pixel using alpha compositing, similar to color rendering: $D=\sum_{i=1}^{M}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value of the i-th Gaussian, $T_{i}$ is the accumulated transmittance, and $\alpha_{i}$ is the opacity value. To suppress floating artifacts while preserving sharp depth discontinuities at object boundaries, we apply anisotropic total variation (TV) regularization to the rendered depth map: $\mathcal{L}_{\text{depth}}=\text{mean}(|\nabla_{x}D|)+\text{mean}(|\nabla_{y}D|)$, where $\nabla_{x}$ and $\nabla_{y}$ are spatial gradients in $x$ and $y$ directions respectively.
@@ 3.5 Masked Gaussian Splatting Optimization
$M_{i}(x,y)=\begin{cases}1&\text{if $(x,y)$ is in an occluded area},\\ 0&\text{if $(x,y)$ is in a static area},\end{cases}$ (12)
where $(x,y)$ represents pixel coordinates in the image. We apply binary dilation to $M_{i}$ for $N_{e}$ iterations, yielding $M^{*}_{i}$. This operation creates a buffer zone around detected dynamic objects, improving the robustness of our static scene reconstruction. The final loss for 3DGS is:
@@
where $I_{i}$, $\hat{I_{i}}$ are reference images and their reconstructions, $\odot$ is the Hadamard product, $\|\cdot\|_{1}$ is L1 norm, $L_{SSIM}$ is a structural similarity loss, $\overline{M}^{*}_{i}$ is a negation of $M^{*}_{i}$ that represents a static background and $\lambda_{\text{SSIM}}$, $\lambda_{\text{L1}}$ and $\lambda_{\text{depth}}$ are weighting factors.
This formulation allows the model to focus on static scene elements, effectively handling dynamic objects in the reconstruction process. By integrating these steps, our method reconstructs static scenes robustly from unconstrained videos while effectively handling transient distractors.